# Variance-Reducing Couplings for Random Features: Perspectives from Optimal Transport

### v1

@@ Appendix A Solving the OT problem for RFFs and RLFs
Denote by $F_{\chi_{d}}(\cdot)$ the cumulative distribution function of $\chi_{d}$. Considering $m=2$ orthogonal frequencies with norms $(\omega_{1},\omega_{2})$, the OT problem in Eq. 6 is solved by the negative monotone coupling
$F_{\chi_{d}}(\omega_{1})+F_{\chi_{d}}(\omega_{2})=1$ (21)
when $\frac{z^{2}}{d}\to 0$ for RFFs and in full generality for RLFs.
@@
Let $A=[0,x_{0}]\times[y_{0},\infty)$, $B=[0,x_{0})\times[0,y_{0})$, $C=[x_{0},\infty)\times[0,y_{0}]$ and $D=(x_{0},\infty]\times(y_{0},\infty]$. Note that $A\cup B\cup C\cup D=\mathbb{R}^{+}\times\mathbb{R}^{+}$ so $\mu^{*}(A\cup B\cup C\cup D)=1$ since the measure is normalised. The subsets are also disjoint apart from $A\cap C=(x_{0},y_{0})$, a singleton of zero measure.222e.g. since $(x_{0},y_{0})\subset\{x_{0}\}\times\mathbb{R}^{+}$ and $\mu^{+}(\{x_{0}\}\times\mathbb{R}^{+})=p_{\chi}(x=x_{0})=0$ since the marginal measure $\chi$ is nonatomic. Eq. 26 implies that $\mu^{*}(B)=0=\mu^{*}(D)$, whereupon $1=\mu^{*}(A\cup C)=\mu^{*}(A)+\mu^{*}(C)=\mu^{*}(A\cup B)+\mu^{*}(C\cup B)$. Now $A\cup B=([0,x_{0}]\times\mathbb{R}^{+})\backslash(\{x_{0}\}\times[0,y_{0}))$ and the set $\{x_{0}\}\times[0,y_{0})$ is zero measure. Therefore, $\mu^{*}(A\cup B)=\mu^{*}([0,x_{0}]\times\mathbb{R}^{+})=F_{\chi_{d}}(x_{0})$. Likewise, $\mu^{*}(C\cup B)=\mu^{*}(\mathbb{R}^{+}\times[0,y_{0}])=F_{\chi_{d}}(y_{0})$. Relabelling $(x_{0},y_{0})\in\Gamma$ by $(\omega_{1},\omega_{2})$, Eq. 21 immediately follows. This completes the proof that negative monotone coupling minimises the kernel estimator variance for orthogonal RLFs.
@@
Let us now turn to the case of RFFs. The optimal transport plan $\mu^{*}$ will instead be $c_{\textrm{RFF}}$-monotone. Unfortunately, Eq. 25 does not hold in general for $c_{\textrm{RFF}}$, so we cannot immediately use the same arguments to conclude that the OT plan is negative monotone. However, Eq. 25 does hold if we just consider the first nontrivial term in the Taylor expansion in $z$ or $\frac{1}{d}$. In particular, dropping terms that are not modified by a coupling (i.e. depend only on $\omega_{1}$ or $\omega_{2}$),
@@
$c_{\textrm{RFF}}(\omega_{1},\omega_{2})=\frac{z^{4}}{\Gamma(\frac{d}{2}+2)}\left[\frac{\omega_{1}^{2}\omega_{2}^{2}}{16}+\mathcal{O}\left(\frac{z^{2}}{d}\right)\right].$ (27)
It is straightforward to show that, if we take the $\frac{z^{2}}{d}\to 0$ limit so we can neglect higher-order terms, $c_{\textrm{RFF}}$ indeed satisfies Eq. 25 and we can make the same arguments about $\Gamma_{\textrm{RFF}}$ as with $\Gamma_{\textrm{RLF}}$. In this limit, the optimal transport plan is again negative monotone. The proof is then complete. ∎
@@
Comments on the $z^{2}/d\to 0$ limit for RFFs. We have seen that for RFFs it is only possible to solve the OT problem in Eq. 6 exactly in the $z^{2}/d\to 0$ asymptotic limit. Here, we comment on why this result is nonetheless interesting and useful. Firstly, note that in practice pairwise norm coupling still substantially suppresses kernel estimator variance, even at data lengthscales chosen by training an exact GP independent of the RF construction (Table 1). Even if at bigger $z$ the method no longer provides the smallest possible estimator variance, it can still substantially reduce it compared to an i.i.d. coupling. Second, from a more theoretical perspective, the $d\to\infty$, small sample regime is exactly where standard QMC methods often fail. It is interesting that OT-driven methods can still provide theoretical guarantees in this low-sample, high-dimensionality setting. Last, we note that, for RFF variance reduction schemes, it is very common to only guarantee gains in the asymptotic limit. This is also the case e.g. for orthogonality [Reid et al., 2023, Yu et al., 2016]: a well-established and widely-used algorithm.
@@
Proof of Corollary 3.4. Supposing that we have $m=d$ frequencies, the sum $\sum_{i,j\neq i}^{d}$ has $d(d-1)$ terms in total. Of these, $2\lfloor\frac{d}{2}\rfloor$ correspond are negative monotone norm couplings, and the remainder are independent. The independent terms are the same in the pairwise norm-coupled and fully i.i.d. configurations, so can be ignored. By Thm. 3.2, we have seen that negative monotone coupling exactly solves the variance reduction OT problem (under mild asymptotic assumptions for RFFs), so these variance contributions will be strictly smaller in the norm-coupled case. It immediately follows that pairwise norm-coupled RFs give strictly lower kernel estimator variance than orthogonal independent-norm RFs, so the result follows. ∎
@@
$\mu^{*}=\arg\inf_{\mu\in\Lambda_{2}(\mathcal{N})}\left[\mathbb{E}_{\bm{\omega}_{1},\bm{\omega}_{2}\sim\mu}\sum_{k=0}^{\infty}\frac{v^{2k}\|\bm{\omega}_{1}+\bm{\omega}_{2}\|_{2}^{k}}{2^{2k}k!\Gamma(k+\frac{d}{2})}\right].$ (28)
@@ A.5 Copulas as numerical OT solvers
[…]
copula_schematic
Figure 3: Schematic of a $2$-dimensional Gaussian copula. Random variables drawn from a Gaussian distribution with unit diagonal covariances (i.e. a correlation matrix) will be marginally normally distributed, $\mathcal{N}(0,1)$. Correlations between the variables will be controlled by the off-diagonal covariances. We can push each coordinate forward using the Gaussian CDF $F_{\mathcal{N}}$ to get marginally uniform random variables, then again with $F_{\chi_{d}}^{-1}$ to get marginal Chi distributions. In Sec. 3.2 and App. A.5, we describe techniques to learn the optimal covariance matrix by gradient descent, directly minimising the kernel approximation error. In doing so, we find a numerical approximation of the solution to the multi-marginal OT problem in Eq. 6.
In the main text, we noted that finding an analytic solution to the multi-marginal OT problem for RFFs and RLFs (Eq. 6) is an open problem. In Sec. 3.2, we briefly presented an alternative numerical approach using copulas. Here, we discuss this in greater detail. Fig. 3 gives a visual overview.
@@
Let $\{\omega_{i}\}_{i=1}^{m}\sim\mathcal{N}(0,\bm{\Sigma})$ where $\bm{\Sigma}\in\mathbb{R}^{m\times m}$ is a correlation matrix, i.e. a positive definite matrix with unit diagonals, and let $F_{\mathcal{N}}$ be the CDF of the standard univariate Gaussian. We say $\{F_{\mathcal{N}}(\omega_{i})\}_{i=1}^{m}$ is distributed according to a Gaussian copula with covariance $\bm{\Sigma},$ and use the notation $\{F_{\mathcal{N}}(\omega_{i})\}_{i=1}^{m}\sim\text{GC}(\bm{\Sigma})$ to denote this.
@@
where $s_{i}=\sqrt{\sum_{j=1}^{i}{\theta_{ij}^{2}}}.$ Note that, since we are directly parameterising the Cholesky factor, we can sample from the associated Gaussian copula with an $\mathcal{O}(m^{2})$ computational cost.
@@
$\mathcal{L}(\bm{\theta})=\mathbb{E}_{\omega_{i}}\left[\sqrt{\frac{1}{N^{2}}\sum_{i,j=1}^{N}(\phi_{\text{RF}}(\bm{x}_{i})^{\top}\phi_{\text{RF}}(\bm{x}_{j})-k(\bm{x}_{i},\bm{x}_{j}))^{2}}~{}\right],$ (30)
where $\omega_{i}\sim\text{GC}(\mathbf{L}_{\bm{\theta}}\mathbf{L}_{\bm{\theta}}^{\top}).$ Note that $\phi_{\text{RF}}$ here depends on $\omega_{i},$ though we have suppressed this dependence for notational simplicity. Assuming that $\phi_{\text{RF}}$ is differentiable with respect to $\omega_{i},$ which is the case in RFFs and RLFs, we can optimise the copula parameters $\bm{\theta}$ by estimating the loss in Eq. 30, computing its gradients with respect to $\bm{\theta},$ and updating its values accordingly.
@@ Appendix B RFF and RLF experimental details
In this appendix, we supplement the discussion in Sec. 3.3, providing more details of our experimental setup for Gram matrix estimation. We also apply norm-coupled RFs to sparse spectrum Gaussian processes [Lázaro-Gredilla et al., 2010], explaining the surprising closing remarks of Sec. 3 that in this case variance reduction does not help downstream performance.
@@ B.1 Details for RFF and RLF experiments
B.2 Do better kernel approximations give better predictions?
@@
Evaluating posterior approximation quality. Table 2 takes the RFs from Sec. 3.3, where we found that our coupling can substantially improve the quality of kernel estimation. It then reports the KL divergence between the exact predictive posterior and the approximate predictive posteriors computed with RFFs and RLFs, respectively. To be clear, Eqs 34 and 36 can be rewritten
@@
Table 2: KL divergences (nats per datapoint) between exact and approximate GP predictive posterior on held out test sets. Note that we use $d$ features for RFFs and $2d$ features for RLFs. Reported errors are equal to two standard errors, i.e. $98\%$ confidence intervals, computed by averaging across splits. Note that differences between data splits are responsible the big errors; within each split we take enough trials that the standard errors become small. For every individual data split, orthogonal is statistically significantly better than i.i.d. but there is no additional benefit from PW norm-coupled. Results differ between splits because we consider a different subset of the dataset.
@@
The left hand panel confirms that we have achieved our stated objective of variance reduction. Moreover, for all couplings the quality of approximation improves as we introduce more ensembles of size $d$. Reading left to right, the other three panels show: (i) the KL divergence between the exact and approximate GP predictive posteriors (as in Table 2), (ii) the predictive RMSE on a held out test set, and (iii) the KL divergence between the exact and approximate GP priors. In every instance, orthogonality provides a substantial gain but, despite the encouraging results for kernel RMSE, there is no additional benefit from PNC. However, it would be wrong to draw the simplistic conclusion that PNC does not give large enough variance savings to see downstream gains: comparable reductions from orthogonality yield substantial improvements. The problem is more fundamental, relating to how the joint distribution of kernel estimates – beyond just the second moment of its pointwise entries – interacts with nonlinear operations like matrix multiplication and inversion.
A non-exhaustive list of publications with a substantial focus on reducing the variance of RF kernel estimates is given by: Yu et al. [2016], Rowland et al. [2018], Reid et al. [2023], Likhosherstov et al. [2022], Yang et al. [2014a], Le et al. [2013], Bojarski et al. [2017], Choromanski et al. [2017], Lyu [2017], Shen et al. [2017], Dao et al. [2017], Munkhoeva et al. [2018]. This can sometimes be a very effective approach for improving downstream performance; e.g. observe the benefits from $\sigma$-coupled GRFs in Sec. 4. But our findings suggest that the true reasons for any boost in performance are more complex than variance reduction alone. There has been some recognition that other properties of the kernel approximation are important e.g. for kernel ridge regression – in particular, the spectral properties of $\widehat{\mathbf{K}}$ [Choromanski et al., 2018, Avron et al., 2017a] – but we believe this to be underappreciated in the literature. More work is needed to fully identify the right desiderata for RF couplings in machine learning, and more algorithms to optimise couplings to improve them. OT provides an excellent framework for achieving this.

### v2

@@ Appendix A Solving the OT problem for RFFs and RLFs
Denote by $F_{\chi_{d}}(\cdot)$ the cumulative distribution function (CDF) of $\chi_{d}$. Consider $m=2$ orthogonal frequencies with norms $(\omega_{1},\omega_{2})$. For RLFs, the OT problem in Eq. 5 is solved by the negative monotone coupling
$F_{\chi_{d}}(\omega_{1})+F_{\chi_{d}}(\omega_{2})=1.$ (20)
For RFFs, Eq. 7 ensures lower cost than any other coupling, provided $z$ is sufficiently small.
@@
Let $A=[0,x_{0}]\times[y_{0},\infty)$, $B=[0,x_{0})\times[0,y_{0})$, $C=[x_{0},\infty)\times[0,y_{0}]$ and $D=(x_{0},\infty]\times(y_{0},\infty]$. Note that $A\cup B\cup C\cup D=\mathbb{R}^{+}\times\mathbb{R}^{+}$ so $\mu^{*}(A\cup B\cup C\cup D)=1$ since the measure is normalised. The subsets are also disjoint apart from $A\cap C=(x_{0},y_{0})$, a singleton of zero measure.111e.g. since $(x_{0},y_{0})\subset\{x_{0}\}\times\mathbb{R}^{+}$ and $\mu^{+}(\{x_{0}\}\times\mathbb{R}^{+})=p_{\chi}(x=x_{0})=0$ since the marginal measure $\chi$ is nonatomic. Eq. 25 implies that $\mu^{*}(B)=0=\mu^{*}(D)$, whereupon $1=\mu^{*}(A\cup C)=\mu^{*}(A)+\mu^{*}(C)=\mu^{*}(A\cup B)+\mu^{*}(C\cup B)$. Now $A\cup B=([0,x_{0}]\times\mathbb{R}^{+})\backslash(\{x_{0}\}\times[0,y_{0}))$ and the set $\{x_{0}\}\times[0,y_{0})$ is zero measure. Therefore, $\mu^{*}(A\cup B)=\mu^{*}([0,x_{0}]\times\mathbb{R}^{+})=F_{\chi_{d}}(x_{0})$. Likewise, $\mu^{*}(C\cup B)=\mu^{*}(\mathbb{R}^{+}\times[0,y_{0}])=F_{\chi_{d}}(y_{0})$. Relabelling $(x_{0},y_{0})\in\Gamma$ by $(\omega_{1},\omega_{2})$, Eq. 20 immediately follows. This completes the proof that negative monotone coupling minimises the kernel estimator variance for orthogonal RLFs.
@@
Let us now turn to the case of RFFs. The optimal transport plan $\mu^{*}$ will instead be $c_{\textrm{RFF}}$-monotone. Unfortunately, Eq. 24 does not hold in general for $c_{\textrm{RFF}}$, so we cannot immediately use the same arguments to conclude that the OT plan is negative monotone. However, it does hold if we just consider the first few terms of its Taylor expansion in $z$. The following is true.
Lemma A.6 (Negative monotone coupling for RFFs).
Denote by $\mu_{\textrm{NM}}$ the negative monotone coupling. Consider a coupling $\mu^{\prime}\in\Lambda_{2}(\eta)\backslash\left\{\mu_{\textrm{NM}}\right\}$, i.e. any other feasible transport map which is not negative monotone. There exists some constant $\delta(\mu^{\prime})>0$ such that $\mathcal{I}(\mu_{\textrm{NM}})<\mathcal{I}(\mu^{\prime})$ for all $z<\delta$ (where $\mathcal{I}(\mu)$ denotes the expectation of the cost function under $\mu$).
Proof. Recall that, for RFFs, the cost function is of the form
$c_{\textrm{RFF}}(\omega_{1},\omega_{2})=\sum_{k=0}^{\infty}\frac{(-1)^{k}z^{2k}\left(\omega_{1}^{2}+\omega_{2}^{2}\right)^{k}}{2^{2k}k!\Gamma(k+\frac{d}{2})},$ (26)
and that we would like to solve $\mu^{*}=\textrm{arg min}_{\mu\in\Lambda_{2}(\eta)}\left[\mathbb{E}_{{{\omega}_{1,2}}\sim\mu}c_{\textrm{RFF}}(\omega_{1},\omega_{2})\right].$ In general, $\mu^{*}=\mu^{*}(z,d)$. This is a very challenging OT problem; we do not believe that a simple closed-form solution exists. However, we can make progress expanding in $k$. The $k=0$ and $k=1$ terms are trivial since they do not contain $\omega_{1,2}$ cross terms. We have
@@
$\mu^{*}=\textrm{arg min}_{\mu\in\Lambda_{2}(\eta)}\left[\frac{z^{4}}{32\Gamma(2+\frac{d}{2})}\mathbb{E}\left((\omega_{1}^{2}+\omega_{2}^{2})^{2}\right)+\sum_{k=3}^{\infty}\frac{(-1)^{k}z^{2k}}{2^{2k}k!\Gamma(k+\frac{d}{2})}\mathbb{E}\left(\left(\omega_{1}^{2}+\omega_{2}^{2}\right)^{k}\right)\right].$ (27)
We can solve the OT problem exactly for the first term, recovering negative monotone coupling. That is,
$\textrm{arg min}_{\mu\in\Lambda_{2}(\eta)}\left[\frac{z^{4}}{32\Gamma(2+\frac{d}{2})}\mathbb{E}\left((\omega_{1}^{2}+\omega_{2}^{2})^{2}\right)\right]=\mu_{\textrm{NM}}$ (28)
for any $z>0$ and $d$ finite. Note also that
$\textrm{max}_{\mu\in\Lambda_{2}(\eta)}\mathbb{E}\left(\left(\omega_{1}^{2}+\omega_{2}^{2}\right)^{k}\right)=2^{k}\mathbb{E}\left(\omega^{2k}\right)$ (29)
because the coupling that maximises the expectation is positive monotone (this follows from our previous arguments almost automatically). This is evaluated as the $k$th moment of a $\chi^{2}_{d}$ distribution, which is nothing other than $2^{k}\Gamma(\frac{k}{2}+d)/\Gamma(\frac{k}{2})$. This means we can upper bound the magnitude of the $k$th term in the expansion by: $z^{2k}/\left(k!\Gamma(\frac{k}{2})\right)$. So for any coupling (feasible transport plan) we can upper bound the magnitude of the sum on the right by: $g(z)\coloneqq\sum_{k=3}^{\infty}z^{2k}/\left(k!\Gamma(\frac{k}{2})\right)$. This goes to 00 as $z\to 0$ and increases monotonically.
Consider some coupling $\mu^{\prime}\in\Lambda_{2}(\eta)\backslash\left\{\mu_{\textrm{NM}}\right\}$. Since the minimiser is unique (if $\mu^{\prime}$ is not negative monotone, it is not a minimiser), there is a positive constant $c$ such that
$c\coloneqq\mathbb{E}_{\mu^{\prime}}\left(\omega_{1}^{2}+\omega_{2}^{2}\right)^{2}-\mathbb{E}_{\mu_{\textrm{NM}}}\left(\omega_{1}^{2}+\omega_{2}^{2}\right)^{2}>0.$ (30)
But we also have that
$\left|\mathbb{E}_{\mu_{\textrm{NM}}}\sum_{k=3}^{\infty}\frac{(-1)^{k}z^{2k}\left(\omega_{1}^{2}+\omega_{2}^{2}\right)^{k}}{2^{2k}k!\Gamma(k+\frac{d}{2})}-\mathbb{E}_{\mu^{*}}\sum_{k=3}^{\infty}\frac{(-1)^{k}z^{2k}\left(\omega_{1}^{2}+\omega_{2}^{2}\right)^{k}}{2^{2k}k!\Gamma(k+\frac{d}{2})}\right|<2g(z).$ (31)
Therefore, to guarantee that $\mathcal{I}(\mu_{\textrm{NM}})<\mathcal{I}(\mu^{\prime})$, it is sufficient that
$\frac{z^{4}}{32\Gamma(2+\frac{d}{2})}c>2g(z).$ (32)
Since $g(z)/z^{4}$ is also monotonically increasing and evaluates to 00 and $z=0$ (by inspecting its Taylor expansion), it is always possible to solve this to find $z=\delta(\mu^{\prime})$ such that, for $z<\delta(\mu^{\prime})$, $\mathcal{I}(\mu_{\textrm{NM}})<\mathcal{I}(\mu^{\prime})$ is guaranteed. This holds for any dimensionality $d$ and any measure $\mu^{\prime}$ that is not negative monotone. ∎
This shows that the negative monotone coupling gives lower expected cost with RFFs than any other coupling if $z$ is sufficiently small, concluding our proof of Thm. 3.2. ∎
@@
Comments on small $z$ for RFFs. We have seen that for RFFs it is only possible to prove optimality of $\mu_{\textrm{NM}}$ when $z$ is small enough. Here, we comment on why this result is nonetheless interesting and useful. Firstly, note that in practice pairwise norm coupling still substantially suppresses kernel estimator variance, even at data lengthscales chosen by training an exact GP independent of the RF construction (Table 1). Even if at bigger $z$ the method no longer provides the smallest possible estimator variance, it can still substantially reduce it compared to an i.i.d. coupling. Second, from a more theoretical perspective, the large $d$, small sample regime is exactly where standard QMC methods often fail. It is interesting that OT-driven methods can still provide theoretical guarantees in this low-sample, high-dimensionality setting. Last, we note that, for RFF variance reduction schemes, it is very common to only guarantee gains in the asymptotic limit. This is also the case e.g. for orthogonality [Reid et al., 2023, Yu et al., 2016]: a well-established and widely-used algorithm.
@@
Proof of Corollary 3.4. Supposing that we have $m=d$ frequencies, the sum $\sum_{i,j\neq i}^{d}$ has $d(d-1)$ terms in total. Of these, $2\lfloor\frac{d}{2}\rfloor$ correspond are negative monotone norm couplings, and the remainder are independent. The independent terms are the same in the pairwise norm-coupled and fully i.i.d. configurations, so can be ignored. By Thm. 3.2, we have seen that negative monotone coupling exactly solves the variance reduction OT problem for RLFs, so these variance contributions will be strictly smaller in the norm-coupled case. It immediately follows that pairwise norm-coupled RLFs give strictly lower kernel estimator variance than orthogonal independent-norm RLFs. For RFFs, Thm. A.6 shows that negative monotone coupling is better than i.i.d. if $z$ is small enough, so the result again follows. ∎
@@
$\mu^{*}=\arg\min_{\mu\in\Lambda_{2}(\mathcal{N})}\left[\mathbb{E}_{\bm{\omega}_{1},\bm{\omega}_{2}\sim\mu}\sum_{k=0}^{\infty}\frac{v^{2k}\|\bm{\omega}_{1}+\bm{\omega}_{2}\|_{2}^{k}}{2^{2k}k!\Gamma(k+\frac{d}{2})}\right].$ (33)
@@ A.5 Copulas as numerical OT solvers
In the main text, we noted that finding an analytic solution to the multi-marginal OT problem for RFFs and RLFs (Eq. 5) is an open problem. In Sec. 3.2, we briefly presented an alternative numerical approach using copulas. Here, we discuss this in greater detail.
@@
Let $\{g_{i}\}_{i=1}^{m}\sim\mathcal{N}(0,\bm{\Sigma})$ where $\bm{\Sigma}\in\mathbb{R}^{m\times m}$ is a correlation matrix, i.e. a positive definite matrix with unit diagonals, and let $F_{\mathcal{N}}$ be the CDF of the standard univariate Gaussian. We say $\{u_{i}\}_{i=1}^{m}$, where $u_{i}\coloneqq F_{\mathcal{N}}(g_{i})$, is distributed according to a Gaussian copula with covariance $\bm{\Sigma}$. We use the notation $\{u_{i}\}_{i=1}^{m}\sim\text{GC}(\bm{\Sigma})$ to denote this.
@@
where $s_{i}=\sqrt{\sum_{j=1}^{i}{\theta_{ij}^{2}}}.$ Note that, since we are directly parameterising the Cholesky factor, we can sample from the associated Gaussian copula with $\mathcal{O}(m^{2})$ computational cost.
@@
$\mathcal{L}(\bm{\theta})=\mathbb{E}_{\{u_{i}\}}\left[\sqrt{\frac{1}{N^{2}}\sum_{i,j=1}^{N}(\phi_{\text{RF}}(\bm{x}_{i})^{\top}\phi_{\text{RF}}(\bm{x}_{j})-k(\bm{x}_{i},\bm{x}_{j}))^{2}}~{}\right],$ (35)
where $\{u_{i}\}\sim\text{GC}(\mathbf{L}_{\bm{\theta}}\mathbf{L}_{\bm{\theta}}^{\top}).$ Note that $\phi_{\text{RF}}$ here depends on $u_{i}$ by pushing forward with $F_{\chi_{d}}^{-1}$ and using the result as random frequency norms, though we have suppressed this dependence for notational simplicity. Assuming that $\phi_{\text{RF}}$ is differentiable with respect to $\omega_{i},$ which is the case in RFFs and RLFs, we can optimise the copula parameters $\bm{\theta}$ by estimating the loss in Eq. 35, computing its gradients with respect to $\bm{\theta},$ and updating its values accordingly.
@@ Appendix B RFF and RLF experimental details
In this appendix, we supplement the discussion in Sec. 3.3, providing more details of our experimental setup for Gram matrix estimation. We also apply norm-coupled RFs to sparse spectrum Gaussian processes [Lázaro-Gredilla et al., 2010], showing that in this case variance reduction does not help downstream performance. We provide experimental details for the Performer [Choromanski et al., 2020] results in Table 2.
@@ B.1 Details for RFF and RLF experiments
B.2 PNC RFFs for Gaussian processes
@@
Evaluating posterior approximation quality. Table 3 takes the RFs from Sec. 3.3, where we found that our coupling can substantially improve the quality of kernel estimation. It then reports the KL divergence between the exact predictive posterior and the approximate predictive posteriors computed with RFFs and RLFs, respectively. To be clear, Eqs 39 and 41 can be rewritten
@@
Table 3: KL divergences (nats per datapoint) between exact and approximate GP predictive posterior on held out test sets. Note that we use $d$ features for RFFs and $2d$ features for RLFs. Reported errors are equal to two standard errors, i.e. $98\%$ confidence intervals, computed by averaging across splits. Note that differences between data splits are responsible for large errors; within each split we take enough trials that the standard errors become small. For every individual data split, orthogonal is statistically significantly better than i.i.d. but there is no additional benefit from PW norm-coupled. Results differ between splits because we consider a different subset of the dataset.
@@
The left hand panel confirms that we have achieved our stated objective of variance reduction. Moreover, for all couplings the quality of approximation improves as we introduce more ensembles of size $d$. Reading left to right, the other three panels show: (i) the KL divergence between the exact and approximate GP predictive posteriors (as in Table 3), (ii) the predictive RMSE on a held out test set, and (iii) the KL divergence between the exact and approximate GP priors. In every instance, orthogonality provides a substantial gain but, despite the encouraging results for kernel RMSE, there is no additional benefit from PNC. However, it would be wrong to draw the simplistic conclusion that PNC does not give large enough variance savings to see downstream gains: comparable reductions from orthogonality yield substantial improvements. The problem is more fundamental, relating to how the joint distribution of kernel estimates – beyond just the second moment of its pointwise entries – interacts with nonlinear operations like matrix multiplication and inversion. Table 4 is a companion to Table 3, reporting predictive RMSEs as in Fig. 6 for the other datasets.
Fourier Features Concrete Abalone CPU Power Airfoil Boston
i.i.d. $1.000$$\pm 0.027$ $1.000$$\pm 0.043$ $1.000$$\pm 0.179$ $1.000$$\pm 0.016$ $1.000$$\pm 0.025$ $1.000$$\pm 0.067$
Halton $1.013$$\pm 0.027$ $1.001$$\pm 0.043$ $1.005$$\pm 0.179$ $0.991$$\pm 0.016$ $0.999$$\pm 0.025$ $1.021$$\pm 0.068$
Sobol $1.010$$\pm 0.025$ $1.000$$\pm 0.045$ $1.004$$\pm 0.138$ $0.991$$\pm 0.022$ $0.997$$\pm 0.029$ $1.027$$\pm 0.058$
Orthogonal $0.917$$\pm 0.026$ $0.994$$\pm 0.044$ $1.019$$\pm 0.182$ $0.915$$\pm 0.019$ $0.930$$\pm 0.025$ $0.974$$\pm 0.066$
+ PNC $0.917$$\pm 0.026$ $0.994$$\pm 0.044$ $1.033$$\pm 0.180$ $0.915$$\pm 0.019$ $0.930$$\pm 0.025$ $0.975$$\pm 0.066$
Laplace Features Concrete Abalone CPU Power Airfoil Boston
i.i.d. $1.000$$\pm 0.028$ $1.000$$\pm 0.043$ $1.000$$\pm 0.192$ $1.000$$\pm 0.017$ $1.000$$\pm 0.029$ $1.000$$\pm 0.073$
Halton $1.007$$\pm 0.028$ $0.997$$\pm 0.044$ $1.003$$\pm 0.190$ $0.964$$\pm 0.017$ $0.989$$\pm 0.029$ $1.019$$\pm 0.075$
Sobol $1.007$$\pm 0.025$ $0.997$$\pm 0.045$ $0.998$$\pm 0.176$ $0.964$$\pm 0.022$ $0.990$$\pm 0.039$ $1.017$$\pm 0.071$
Orthogonal $0.979$$\pm 0.027$ $0.990$$\pm 0.044$ $1.003$$\pm 0.187$ $0.899$$\pm 0.018$ $0.958$$\pm 0.028$ $0.979$$\pm 0.071$
+ PNC + antithetic $0.981$$\pm 0.027$ $0.991$$\pm 0.044$ $1.006$$\pm 0.187$ $0.909$ ±$\pm 0.018$ $0.958$$\pm 0.028$ $0.982$$\pm 0.071$
Table 4: Predictive RMSEs (normalised to RMSE of i.i.d.) of the approximate GP predictive posteriors on held out test sets. Reported errors are equal to two standard errors, i.e. $98\%$ confidence intervals, computed by averaging across splits. Substantially lower kernel estimator variance does not guarantee better predictive performance, though performance tends to be no worse.